OpenStudy (anonymous):

The area of a trapezoid is given by the formula A 1/2= (b 1 + b 2)h, where base b 1 is parallel to base b 2 and h is the height. Solve the formula for h

OpenStudy (paxpolaris):

\[\large A=\frac12\left( b_1+b_2 \right)\cdot h\]you want to re-write this equation so that you only left with \(h\) on one side of the equal to sign.

jhonyy9 (jhonyy9):

1 A = --- ( b_1 +b_2)*h 2 2A = (b_1+b_2)*h 2A h = ----------- (b_1 +b_2) hope this will help you